# Metric Oja depth, new statistical tool for estimating the most central objects

### v2

@@ Abstract
The Oja depth (simplicial volume depth) is one of the classical statistical techniques for measuring the central tendency of data in multivariate space. Despite the widespread emergence of object data like images, texts, matrices or graphs, a well-developed and suitable version of Oja depth for object data is lacking. To address this shortcoming, a novel measure of statistical depth, the metric Oja depth applicable to any object data, is proposed. Two competing strategies are used for optimizing metric depth functions, i.e., finding the deepest objects with respect to them. The performance of the metric Oja depth is compared with three other depth functions (half-space, lens, and spatial) in diverse data scenarios.
@@ 1.1 Background
Nowadays, data is being generated in high volumes, at high speeds, and with great diversity. A key aspect of modern data analysis is that we no longer deal exclusively with data existing in Euclidean spaces. Instead, data are currently taking on more complex formats, such as images, graphs, matrices, etc., all of which reside in non-Euclidean spaces and are collectively known as object data. Methodology for object data, known as object data analysis or metric statistics [8], are currently bountiful in statistical literature, see, e.g., [9, 29, 30, 31]. Object data are typically modelled as residing in an arbitrary metric space $(\mathcal{X},d)$ and we also adopt this approach.
@@
The foundation for depth functions was introduced in 1975 by Tukey who proposed his seminal half-space depth for multivariate data to rank observations and reveal features of the underlying data distribution [25]. Apart from this, numerous depth functions for data in Euclidean space have been proposed. E.g., the convex hull peeling depth (onion depth) [1, 12], simplicial volume depth (Oja depth) [22] and several others, see Table A.1 in Appendix A and the comprehensive addressing in [19].
@@
In recent years, object/metric versions of various depth functions $D$ have been proposed, see Table A.1. That is, given an object $X\in\mathcal{X}$ and a distribution $P$ taking values in $\mathcal{X}$, the depth $D(X;P)$ describes how central the object $X$ is w.r.t. $P$, much in the same way as for the Euclidean depths earlier. Two key properties of these extensions are that (a) they depend on the data only through the metric $d$, making them applicable to any form of object data, regardless in which specific metric space they live, and (b) if $(\mathcal{X},d)$ is a Euclidean space, then the original Euclidean version of the depth is recovered, showing that these metric depth functions are “true” generalizations of their classical counterparts. Metric lens depth was developed in [16, 5, 14], metric half-space depth was proposed in [7] and metric spatial depth in [27].
@@
Most of the depth functions in Table A.1, in particular all three that have object versions, are robust [18], meaning that their performance is not skewed by the presence of possible outliers. In Section 2 we list the robustness properties of the existing metric depth functions, and more details on the robustness of Euclidean depths can be found in [21, Table 1] and [32, 20]. Note that not all location estimation procedures are robust. For example, the Fréchet mean [13] which is a generalization of the concept of average and a classical estimator of location for object data, is not robust, but is instead affected greatly by data outliers. Despite the popularity of Fréchet mean and similar tools, in this work we concentrate solely on robust methods. This is because object data is very diverse and recognizing any possible outliers in them can, as such, be very difficult. To ensure that our analyses stay reliable, we therefore find it imperative to develop robust methods, which automatically protect us against outliers, regardless of their exact type and whether we recognize them or not.
@@ 1.2 Our Contributions
(i) We propose a new depth function for object data, the metric Oja depth, and extensively study its theoretical properties, in particular characterizing its range of values and showing the sample consistency of its maximizers under specific assumptions.
@@ 1.3 Contents
We review the three existing metric depth functions in Section 2. In Section 3, we introduce the proposed new metric depth function and develop its theoretical properties. In Section 4, we evaluate and compare it with three other depth functions, using two simulation scenarios. In Section 5, we further test our metric depth function on a real dataset, relying on computational techniques to conduct statistical inference. Finally, future recommendations and conclusions are compiled in Section 6.
@@ 2 Review of existing metric depth functions
Every pair of objects $x_{1},x_{2}$ in the infimum divides the space $\mathcal{H}$ in two subsets (objects closer to $x_{1}$ than $x_{2}$ and vice versa). The depth $D_{H}(x)$ is defined to be the smallest possible probability mass of a subset produced in this way and containing the object $x$. As with the metric lens depth, it is easy to see that, for an outlying object $x$, it is easy to find a subset which contains $x$ but has very little $P$-probability mass, making $D_{H}(x)$ small. An approximative algorithm, that we also use in this paper to compute the metric half-space depth, is given in [7, Algorithm 1].
@@
Throughout this work, we refer to metric lens depth, metric half-space depth and metric spatial depth as MLD, MHD and MSD, respectively. All of these have the property of vanishing at infinity meaning that, for objects far enough away from the bulk of the distribution $P$, the depth values approach zero. All three also satisfy specific forms of continuity (small changes in $x$ entail small changes in the depth) and both MLD and MHD are invariant to data transformations that preserve the ordering of distances. Regarding robustness properties, MHD and MSD are robust, the former in the sense of having a high breakdown point and the latter in the sense of having a bounded influence function. Whereas, MLD is also likely to be highly robust, being based only on comparisons between distances and not their absolute sizes, but this aspect of MLD has not yet been studied in the literature. For detailed statements of these, and some other properties of the three depth functions, including sample consistency results, we refer the reader to the papers [5, 14, 7, 27].
@@ 3.1 Population-level properties
$D_{O3}(x):=\frac{1}{1+\mathrm{E}\left[\left\{|B_{3}(x,X_{1},X_{2},X_{3})|+4d^{2}(x,X_{1})d^{2}(x,X_{2})d^{2}(x,X_{3})\right\}^{1/2}\right]}.$ (2)
The subscript O refers to “Oja” due to the close connection between $D_{O3}(x)$ and the classical Oja depth explained in Appendix B. By Theorem C.1 in Appendix C, taking the square root in the denominator of (2) is well-defined. The intuitive meaning of $D_{O3}(x)$ is not clear from its definition (2) alone, so we next present a collection of results that allow better understanding it, starting with a moment condition that guarantees its existence.
@@
Assume that, for some $a\in\mathcal{X}$, we have $\mathrm{E}\{d(X,a)\}<\infty$. Then $D_{O3}(x)$ exists as well-defined.
By “well-defined” in Theorem 1 we mean that the expected value used to compute $D_{O3}(x)$ exists as finite. The condition $\mathrm{E}\{d(X,a)\}<\infty$ is analogous to requiring a univariate random variable to have a finite mean and is a quite mild condition as even the Frechét mean already requires the existence of second moments. The choice of $a$ in Theorem 1 is completely arbitrary as, if the condition $\mathrm{E}\{d(X,a)\}<\infty$ holds for some $a$, then, by the triangle inequality, it holds for all $a\in\mathcal{X}$.
@@
Theorem 2 implies that large values of $D_{03}(x)$ are indicative of the centrality of the object $x$ with respect to the distribution $P$. That is, under the extremal case $D_{03}(x)=1$ the object $x$ must almost surely lie in between of at least two of three randomly sampled objects $X_{1},X_{2},X_{3}$ from $P$.
@@
Assume that, for some $a\in\mathcal{X}$, we have $\mathrm{E}\{d^{2}(X,a)\}<\infty$. Let $x_{n}$ be a divergent sequence of points in $\mathcal{X}$. Then, $D_{O3}(x_{n})\rightarrow 0$ as $n\rightarrow\infty$.
The second moment condition in Theorem 3 is used to avoid certain pathological behavior, see the proof of the result in Appendix C. Theorem 3 states that small values of $D_{O3}(x)$ are reached by outlying objects, i.e., points that are located far away from the bulk of the distribution $P$.
@@ 3.2 Sample-level properties
$D_{O3,n}(x):=\frac{1}{1+\frac{1}{\binom{n}{3}}\sum_{(i,j,k)}h(x,X_{i},X_{j},X_{k})},$
@@
Standard results on $U$-statistics [17] with the continuous mapping theorem and Theorem 1 gives the following point-wise convergence result which states that the depth of any single object $x$ can be accurately estimated using $D_{O3,n}$.
@@
Assume that, for some $a\in\mathcal{X}$, we have $\mathrm{E}\{d(X,a)\}<\infty$. Then, for any fixed $x\in\mathcal{X}$, we have $D_{O3,n}(x)\rightarrow_{p}D_{O3}(x)$ as $n\rightarrow\infty$.
@@
The classical $M$-estimator argument [26, Theorem 5.7] then shows that the point with maximal depth, assuming that it is unique, can be consistently estimated with the sample depth function $D_{O3,n}$.
@@
The sample maximizers in Corollary 6 exist since $D_{O3,n}$ is a continuous function over a totally bounded and complete, and hence compact, set. The assumption of a unique population maximizer is discussed further in Section 6. Also MLD and MHD satisfy analogous consistency properties, see [5, 14, 7].
@@ 3.3 Auxiliary depth-like quantity
Our proposed depth $D_{O3}$ is based on the 3-dimensional Oja depth, see Appendix B for details, and a natural question is whether also the $p$-dimensional Oja depth can be used for similar purposes with $p\neq 3$. In this section we briefly investigate the version with $p=2$. Letting $B_{2}(x_{0},x_{1},x_{2})$ denote the $2\times 2$ top left sub-matrix of $B_{3}(x_{0},x_{1},x_{2},x_{3})$, we define
$D_{O2}(x):=\frac{1}{1+\mathrm{E}\left\{|B_{2}(x,X_{1},X_{2})|^{1/2}\right\}},$ (3)
which, as shown in Appendix B, is a metric generalization of the classical 2-dimensional Oja depth. As with $D_{O3}$, also $D_{O2}$ is well-defined as soon as the first moment exists, i.e., $\mathrm{E}\{d(X,a)\}<\infty$ for some $a\in\mathcal{X}$ (proof omitted). However, the next result, which follows directly from Theorem C.1(i) in Appendix C shows that $D_{O2}(x)$ does not actually measure the centrality of $x$ w.r.t. $P$ and is thus not a proper depth measure.
@@
Intuitively, $D_{O2}(x)$ takes a large value if, for random $X_{1},X_{2}$ and the fixed object $x$, one of the three is typically in between the two others. This does not yet guarantee that $x$ is central with respect to $P$ as it is possible that $X_{1}$ (or $X_{2}$) is the in-between point. As such, large values of $D_{O2}(x)$ do not characterize centrality. As an extreme example of this, if $(\mathcal{X},d)$ is the one-dimensional Euclidean space, then $D_{O2}(x)=1$ for all points $x\in\mathbb{R}$. Due to this, we do not pursue the theoretical properties of $D_{O2}$ further. However, we still include $D_{O2}(x)$ in our data examples, where it works surprisingly well in locating the centers of distributions.

### v3

@@ Abstract
The Oja depth (simplicial volume depth) is one of the classical statistical techniques for measuring the central tendency of data in multivariate space. Despite the widespread emergence of object data like images, texts, matrices or graphs, a well-developed and suitable version of Oja depth for object data is lacking. To address this shortcoming, a novel measure of statistical depth, the metric Oja depth applicable to any object data, is proposed. Two competing strategies are used for optimizing metric depth functions, i.e., for finding the deepest objects with respect to them. The performance of the metric Oja depth is compared with three other depth functions (half-space, lens, and spatial) in diverse data scenarios.
@@ 1.1 Background
Nowadays, data are being generated in high volumes, at high speeds, and with great diversity. A key aspect of modern data analysis is that we no longer deal exclusively with data existing in Euclidean spaces. Instead, data are currently taking on more complex formats, such as images, graphs, matrices, etc., all of which reside in non-Euclidean spaces and are collectively known as object data. Methodology for object data, known as object data analysis or metric statistics [9], are currently abundant in statistical literature; see, e.g., [10, 37, 41, 42]. Object data are typically modelled as residing in an arbitrary metric space $(\mathcal{X},d)$ and we also adopt this approach.
@@
The foundation for depth functions was introduced in 1975 by Tukey who proposed his seminal half-space depth for multivariate data to rank observations and reveal features of the underlying data distribution [34]. Apart from this, numerous depth functions for data in Euclidean space have been proposed. E.g., the convex hull peeling depth (onion depth) [1, 13], simplicial volume depth (Oja depth) [30] and several others; see Table 1 and the comprehensive addressing in [27].
Author Year Depth function Has a metric version Complexity
Mahalanobis [25] 1936 Mahalanobis depth (distance) No -
Tukey [34] 1975 Half-space/location/Tukey depth Yes [7] $\mathcal{O}(n^{2}h)+\mathcal{O}(n^{3})$
Barnett, Eddy [1, 13] 1976, 1981 Convex hull peeling/onion depth No -
Liu; Zuo & Serfling [21, 43] 1992, 2000 Projection depth No -
Liu [20] 1990 Simplicial depth No -
Oja, Zuo & Serfling [30] 1983, 2000 Simplicial volume depth/Oja depth Yes, in the current paper $\mathcal{O}(n^{2}h)+\mathcal{O}(n^{4})$
Koshevoy & Mosler [18] 1997 Zonoid depth No -
Vardi & Zhang, Serfling [36, 31] 2000, 2002 Spatial depth Yes [39] $\mathcal{O}(n^{2}h)+\mathcal{O}(n^{3})$
Liu & Modarres [22] 2011 Lens depth Yes [17, 6, 15] $\mathcal{O}(n^{2}h)+\mathcal{O}(n^{3})$
Yang & Modarres [40] 2018 $\beta$-skeleton depths Not for general $\beta$, see the caption -
Table 1: List of depth functions. In the bottom row, the metric $\beta$-skeleton depth is considered only for $\beta=2$, where the concept reduces to lens depth. The listed complexities are for computing depths on the full training sample of size $n$; For metric half-space depth, the complexity assumes the use of [7, Algorithm 1] with the sample itself as anchors. The complexities consist of two parts, distance matrix formation and depth computation, where in the former, $h$ denotes the cost of computing the distance between a single pair of objects in $(\mathcal{X},d)$; see Section 3.2 for further details.
@@
In recent years, object/metric versions of various depth functions $D$ have been proposed; see the second-to-last column of Table 1. That is, given an object $X\in\mathcal{X}$ and a distribution $P$ taking values in $\mathcal{X}$, the depth $D(X;P)$ describes how central the object $X$ is w.r.t. $P$, much in the same way as for the Euclidean depths earlier. Two key properties of these extensions are that (a) they depend on the data only through the metric $d$, making them applicable to any form of object data, regardless in which specific metric space they live, and (b) if $(\mathcal{X},d)$ is a Euclidean space, then the original Euclidean version of the depth is recovered, showing that these metric depth functions are “true” generalizations of their classical counterparts. Metric lens depth was developed in [17, 6, 15], metric half-space depth was proposed in [7] and metric spatial depth in [39].
@@
Most of the depth functions in Table 1, in particular all three that have object versions, are robust [26], meaning that their performance is not skewed by the presence of possible outliers. In Section 2 we list the robustness properties of the existing metric depth functions, and more details on the robustness of Euclidean depths can be found in [28, Table 1] and [43, 29]. Note that not all location estimation procedures are robust. For example, the Fréchet mean [14], which is a generalization of the concept of average and a classical estimator of location for object data, is not robust, but is instead affected greatly by data outliers. Despite the popularity of Fréchet mean and similar tools, in this work we concentrate solely on robust methods. This is because object data are very diverse, which makes recognizing any possible outliers very difficult. Robust analysis methods protect against the influence of outliers and leads to reliable results, regardless of the exact type of outliers and whether they are detected in the data or not.
@@ 1.2 Our Contributions
(i) We propose a new depth function for object data, the metric Oja depth, and extensively study its theoretical properties. In particular, we characterize its range of values and show the sample consistency of its maximizers under specific assumptions.
@@ 1.3 Contents
We review the three existing metric depth functions in Section 2. In Section 3, we introduce the proposed new metric depth function and develop its theoretical properties. In Section 4, we evaluate and compare it with three other depth functions, using two simulation scenarios. In Section 5, we further test our metric depth function on a real dataset, using permutation and rank tests to obtain inferential results. Finally, future recommendations are compiled in Section 6.
@@ 2 Review of existing metric depth functions
Every pair of objects $x_{1},x_{2}$ in the infimum divides the space $\mathcal{H}$ in two subsets (objects closer to $x_{1}$ than $x_{2}$ and vice versa). The depth $D_{H}(x)$ is defined to be the smallest possible probability mass of a subset produced in this way and containing the object $x$. As with the metric lens depth, it is easy to see that, for an outlying object $x$, it is easy to find a subset which contains $x$ but has very small $P$-probability mass, making $D_{H}(x)$ small. An approximative algorithm, that we also use in this paper to compute the metric half-space depth, is given in [7, Algorithm 1].
@@
$D_{S}(x)=1-\frac{1}{2}\mathbb{E}\left[\mathbb{I}(d(X_{1},x)\neq 0,d(X_{2},x)\neq 0)\left\{\frac{d^{2}(X_{1},x)+d^{2}(X_{2},x)-d^{2}(X_{1},X_{2})}{d(X_{1},x)d(X_{2},x)}\right\}\right],$
@@
Throughout this work, we refer to metric lens depth, metric half-space depth and metric spatial depth as MLD, MHD and MSD, respectively. All of these have the property of vanishing at infinity meaning that, for objects far enough away from the bulk of the distribution $P$, the depth values approach zero. All three also satisfy specific forms of continuity (small changes in $x$ cause small changes in the depth) and both MLD and MHD are invariant to data transformations that preserve the ordering of distances. Regarding robustness properties, MHD and MSD are robust, the former in the sense of having a high breakdown point and the latter in the sense of having a bounded influence function. Whereas, MLD is also likely to be highly robust, being based only on comparisons between distances and not their absolute sizes, but this aspect of MLD has not yet been studied in the literature. For detailed statements of these, and some other properties of the three depth functions, including sample consistency results, we refer the reader to the papers [6, 15, 7, 39].
@@ 3.1 Population-level properties
$D_{O3}(x):=\frac{1}{1+\mathbb{E}\left[\left\{|B_{3}(x,X_{1},X_{2},X_{3})|+4d^{2}(x,X_{1})d^{2}(x,X_{2})d^{2}(x,X_{3})\right\}^{1/2}\right]}.$ (2)
The subscript O refers to “Oja” due to the close connection between $D_{O3}(x)$ and the classical Oja depth explained in Appendix A. By Theorem B.1 in Appendix B, taking the square root in the denominator of (2) is well-defined. The intuitive meaning of $D_{O3}(x)$ is not clear from its definition (2) alone, so we next present a collection of results that allow better understanding it, starting with a moment condition that guarantees its existence.
@@
Assume that, for some $a\in\mathcal{X}$, we have $\mathbb{E}\{d(X,a)\}<\infty$. Then $D_{O3}(x)$ exists as well-defined.
By “well-defined” in Theorem 1 we mean that the expected value used to compute $D_{O3}(x)$ exists as finite. The condition $\mathbb{E}\{d(X,a)\}<\infty$ is analogous to requiring a univariate random variable to have a finite mean and is a rather mild condition as even the Frechét mean already requires the existence of second moments. The choice of $a$ in Theorem 1 is completely arbitrary as, if the condition $\mathbb{E}\{d(X,a)\}<\infty$ holds for some $a$, then, by the triangle inequality, it holds for all $a\in\mathcal{X}$.
@@
Theorem 2 implies that large values of $D_{O3}(x)$ are indicative of the centrality of the object $x$ with respect to the distribution $P$. That is, under the extremal case $D_{O3}(x)=1$ the object $x$ must almost surely lie in between of at least two of three randomly sampled objects $X_{1},X_{2},X_{3}$ from $P$.
@@
Assume that, for some $a\in\mathcal{X}$, we have $\mathbb{E}\{d^{2}(X,a)\}<\infty$. Let $x_{n}$ be a divergent sequence of objects in $\mathcal{X}$. Then, $D_{O3}(x_{n})\rightarrow 0$ as $n\rightarrow\infty$.
The second moment condition in Theorem 3 is used to avoid certain pathological behavior; see the proof of the result in Appendix B. Theorem 3 states that small values of $D_{O3}(x)$ are reached by outlying objects, i.e., points that are located far away from the bulk of the distribution $P$.
@@ 3.2 Sample-level properties
$D_{O3,n}(x):=\frac{1}{1+\frac{1}{\binom{n}{3}}\sum_{(i,j,k)}h(x,X_{i},X_{j},X_{k})},$ (3)
@@
Standard results on $U$-statistics [19] combined with the continuous mapping theorem and Theorem 1 give the following point-wise convergence result which states that the depth of any single object $x$ can be accurately estimated using $D_{O3,n}$.
@@
Assume that, for some $a\in\mathcal{X}$, we have $\mathbb{E}\{d(X,a)\}<\infty$. Then, for any fixed $x\in\mathcal{X}$, we have $D_{O3,n}(x)\rightarrow_{p}D_{O3}(x)$ as $n\rightarrow\infty$.
@@
The classical $M$-estimator argument [35, Theorem 5.7] then shows that the point with maximal depth, assuming that it is unique, can be consistently estimated with the sample depth function $D_{O3,n}$.
@@
The sample maximizers in Corollary 6 exist since $D_{O3,n}$ is a continuous function on a totally bounded and complete, and hence compact, set. The assumption of a unique population maximizer is discussed further in Section 6. Also MLD and MHD satisfy analogous consistency properties; see [6, 15, 7].
We conclude the subsection by discussing computational aspects of the proposed $D_{O3,n}$. Obtaining the depths of all sample points comprises of (A) computing the distance matrix $\{d(X_{i},X_{j})\}_{i,j=1}^{n}$ and, (B) using this to compute the depths of the sample points with (3). The cost of step (A), distance computation, depends on the metric space in question and is essentially $O(n^{2}h)$ where $h$ is the cost of computing the distance between a single pair of objects in $(\mathcal{X},d)$. For example, both in $\mathbb{R}^{p}$ with the Euclidean metric and on the unit sphere in $\mathbb{R}^{p}$ with the arc length metric, forming the distance matrix is an $O(n^{2}p)$-operation, whereas on the manifold of $p\times p$ positive definite matrices with either the log-Euclidean or Riemannian metric, this operation is $O(n^{2}p^{3})$. Importantly, step (A) is identical for all metric depth functions, implying that their complexities can be compared based on step (B) only, detailed next.
Step (B), computing the actual depths, is done based on the $n\times n$ distance matrix and, as such, the complexity of this step depends only on $n$ and not on other aspects of the sample. In particular, the dimensionality of the data plays no role in this step. The computational complexity of our proposed depth is from (3) seen to be $\mathcal{O}(n^{4})$, whereas the complexity of MLD, MSD and MHD is $\mathcal{O}(n^{3})$, for computing the depths of the full sample (assuming [7, Algorithm 1] with the sample itself as anchors is used for MHD). The time efficiency of our proposed depth could be improved, at the cost of accuracy, by using partial U-statistics to randomly choose which elements to include in the triple sum in (3). However, we have not resorted to this compromise here. The previous information on the complexities has been summarized also in the final column of Table 1. The complexities of the Euclidean counterparts of these depths and other Euclidean depths can be found in [29, Table 1.9].
The computation of MLD and MHD does not rely on the actual distances $d(X_{i},X_{j})$ but only on indicators of the form $\mathbb{I}\{d(X_{i},X_{j})\leq d(X_{i},X_{k})\}$; see Section 2. This means that MLD and MHD (a) are invariant to transformations that preserve these indicators, and (b) can be computed even if one knows only the binary values of these triple comparisons. The latter viewpoint for MLD was taken in [17] to find deepest objects based on binary dissimilarity data. Our proposed depth, on the other hand, uses the actual values $d(X_{i},X_{j})$ in its computation, meaning that it cannot be computed based on indicators alone. This limits its usefulness in such dissimilarity scenarios but, at the same time, allows our depth to use additional information from the data, which later manifests as improved performance in several experiments compared with the other methods. This behavior is perfectly in line with the original Oja depth, which also depends on the numerical values of the distances $\|X_{i}-X_{j}\|$.
@@ 3.3 Auxiliary depth-like quantity
Our proposed depth $D_{O3}$ is based on the 3-dimensional Oja depth; see Appendix A for details, and a natural question is whether also the $p$-dimensional Oja depth can be used for similar purposes with $p\neq 3$. In this section we briefly investigate the version with $p=2$. Letting $B_{2}(x_{0},x_{1},x_{2})$ denote the $2\times 2$ top left sub-matrix of $B_{3}(x_{0},x_{1},x_{2},x_{3})$, we define
$D_{O2}(x):=\frac{1}{1+\mathbb{E}\left\{|B_{2}(x,X_{1},X_{2})|^{1/2}\right\}},$ (4)
which, as shown in Appendix A, is a metric generalization of the classical 2-dimensional Oja depth. As with $D_{O3}$, also $D_{O2}$ is well-defined as soon as the first moment exists, i.e., $\mathbb{E}\{d(X,a)\}<\infty$ for some $a\in\mathcal{X}$ (proof omitted). However, the next result, which follows directly from Theorem B.1(i) in Appendix B shows that $D_{O2}(x)$ does not actually measure the centrality of $x$ w.r.t. $P$ and is thus not a proper depth measure.
@@
Intuitively, $D_{O2}(x)$ takes a large value if, for random $X_{1},X_{2}$ and the fixed object $x$, one of the three objects is typically located between the other two. This, however, does not guarantee that $x$ is central with respect to $P$, as it is possible that $X_{1}$ (or $X_{2}$) is the in-between point. Thus, large values of $D_{O2}(x)$ do not characterize centrality. As an extreme example, if $(\mathcal{X},d)$ is the one-dimensional Euclidean space, then $D_{O2}(x)=1$ for all $x\in\mathbb{R}$. Consequently, we do not pursue the theoretical properties of $D_{O2}$ further. However, we still include $D_{O2}(x)$ in our data examples, where it performs surprisingly well in locating the centers of distributions.